A call and put expire in 0.41 year and have an exercise price of $100. The underlying stock is worth $90 and has a standard deviation of 0.25. The annual risk-free rate is 11 percent. The annual dividend yield (q) on the stock is 2%. The put option price from the three-period binomial model is:

8.67

9.467

9.207

7.593
Answer #1

(الم) م * ga da = di- 6xth (Y2) As per Black Scholes Model value of put option = N(da) xkx where the values are Current priceHOPE THIS ANSWER IS USEFUL. PLEASE LIKE THE ANSWER